# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v4

@@ Abstract
Knowledge of the effect of interventions, known as the treatment effect, is paramount for decision-making. Approaches to estimating this treatment effect using conditional average treatment effect (CATE) meta-learners often provide only a point estimate of this treatment effect, while additional uncertainty quantification is frequently desired to enhance decision-making confidence. To address this, we introduce two novel approaches: the conformal convolution T-learner (CCT-learner) and conformal Monte Carlo (CMC) meta-learners. The approaches leverage weighted conformal predictive systems (WCPS), Monte Carlo sampling, and CATE meta-learners to generate predictive distributions of individual treatment effect (ITE) that could enhance individualized decision-making. Although we show how assumptions about the noise distribution of the outcome influence the uncertainty predictions, our experiments demonstrate that the CCT- and CMC meta-learners achieve strong coverage while maintaining narrow interval widths. They also generate probabilistically calibrated predictive distributions, providing reliable ranges of ITEs across various synthetic and semi-synthetic datasets.
Code: https://github.com/predict-idlab/cct-cmc.
@@ 1 Introduction
"To act or not to act?" is a crucial question in decision-making across domains such as healthcare, policy formulation, and marketing [1, 2]. Today, decisions rely on the average results of studies such as clinical trials, A/B tests, or surveys. However, there is an increasing demand for more individualized interventions to optimize outcomes when a decision has to be made [1].
Specialized machine learning (ML) models offer solutions by estimating the conditional average treatment effect (CATE), predicting whether an intervention will be beneficial given the context [3, 4, 5, 6, 7]. However, CATE approaches typically provide single-point estimates, limiting their applicability in uncertain environments, especially in higher-risk settings such as healthcare [8, 9].
Uncertainty quantification of these CATE estimates and variations of individual treatment effect (ITE) addresses these limitations by capturing both the overall uncertainty in the estimate (epistemic) and the potential variation in ITEs (aleatoric). Traditionally, Bayesian approaches, such as Bayesian additive regression trees and Gaussian processes, have been used [10, 11] but they often suffer from under coverage and lack model-agnosticism, limiting their compatibility with modern ML models [12]. Alternatively, frequentist approaches, such as causal forests [3], only quantify the (epistemic) uncertainty around the CATE estimate, whereas a quantile regression approach to potential outcomes using the pinball loss function [13] only gives prediction intervals for the counterfactual outcomes and requires a Bonferroni type of correction, resulting in highly conservative intervals for ITE.
Conformal prediction (CP) [14] offers a promising alternative, providing distribution-free prediction intervals with finite-sample coverage guarantees under the exchangeability assumption. CP can be leveraged to generate prediction intervals with conservative coverage guarantees. Weighted CP accounts for the violation of the exchangeability assumption due to possible covariate shift during counterfactual inference [12]. Alternatively, one can alleviate the covariate shift by applying CP on the pseudo-outcomes of the two-step meta-learner [15]. While both approaches address calibration issues, they are often overly conservative and only generate prediction intervals, offering limited insights compared to predictive distributions which enable to construct decision-making systems by optimizing utility functions [16].
Contributions
This work introduces the conformal convolution T-learner (CCT-learner), a framework built on single-step meta-learners, that leverages convolutions and (weighted) conformal predictive systems (WCPS) [17, 18] to provide predictive distributions of ITE outcomes and less conservative prediction intervals. The CCT-learner outperforms existing prediction interval approaches for ITE by employing WCPS for counterfactual inference, addressing covariate shifts through propensity score-based weighting. The resulting calibrated distributions for treated and control outcomes are then convoluted to obtain the ITE distribution. Although our approach assumes conditionally independent predictive uncertainty of (potential) outcomes, we still have conservatively valid prediction intervals if the predictive uncertainty is positively correlated. Additionally, we argue that conditional independence of the (potential) outcomes aleatoric uncertainty is a needed assumption to provide calibrated predictive distributions and postulate this as the epsilon problem.
We demonstrate the CCT-learner’s effectiveness through simulations on three synthetic and three semi-synthetic datasets. As the convolution operation on nonparametric distributions during inference incurs significant computational cost, we additionally introduce the conformal Monte Carlo (CMC) T-learner that replaces the convolution with Monte Carlo sampling during training. This approach generates pseudo-outcomes for training a CPS, resulting in faster inference times. Simulations show that the CMC T-learner achieves coverage levels comparable to the CCT-learner with minimal efficiency loss. The introduced CMC framework allows adaptations to other meta-learners, such as the S-learner and X-learner [4]. We also establish the existence of a universal CCT- and CMC T-learner which converges to the true data-generating distribution as training data grows infinitely.
2 Background
Problem definition
@@ 2 Background
In this work, we use the Neyman-Rubin [19] potential outcome framework to formulate and discuss our problem. We assume access to a sample of observations $\mathcal{D}=\{(Y_{i},X_{i},W_{i})\}^{n}_{i=1}$, with $(Y_{i},X_{i},W_{i})\overset{\mathrm{iid}}{\sim}\mathcal{P}$ with $\mathcal{P}$ the resulting joint distribution of the data generation process. Here, $Y_{i}\in\mathcal{Y}$ represents the continuous outcome of interest, $X_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ denotes the per-person feature values, and $W_{i}\in\{0,1\}$ is the binary treatment assignment. The joint distribution $\mathcal{P}$ is defined by following the covariate distribution and nuisance functions:
$\pi(x)$ $=\mathbb{P}(W=1|X=x)$ $\mu_{w}(x)$ $=\mathbb{E}(Y|X=x,W=w)$
$X$ $\sim\Lambda,$ $W$ $\sim\text{Bernoulli}(\pi(x))$
$Y^{w}$ $=\mu_{w}(X)+\varepsilon^{w},$ $w$$\in\{0,1\}$
$Y$ $=Y^{0}(1-W)+Y^{1}W$ $\tau(x)$ $=\mathbb{E}(Y^{1}-Y^{0}|X=x)$
$ITE$ $=Y^{1}-Y^{0}=\tau+\varepsilon_{a,ITE}$ $\varepsilon_{a,ITE}$ $=\varepsilon^{1}-\varepsilon^{0}$
@@
where $\pi(x)$ denotes the propensity score that defines the probability of treatment assignment for an individual, $\mu_{w}(x)$ represents the expected (potential) outcome function of an individual for treatment assignment $w$, $\Lambda$ represents the marginal distribution of $X$, $Y^{w}$ represent the outcome when an individual receives treatment $w$, $\varepsilon^{w}$ are zero-mean random variables independent of $X$ and $W$, $Y$ is the observed outcome, $\tau(x)$ denotes the CATE function, $ITE$ the individual treatment effect, and $\varepsilon_{a,ITE}$ represent a zero-mean random variable. Next, we assume that the described data-generating process follows the following assumptions: (1) Consistency: $Y^{w_{i}}=Y_{i}$; (2) Unconfoundedness: $Y^{0},Y^{1}\perp W|X$ or equivalently $\varepsilon(0),\varepsilon(1)\perp W|X$; and (3) Positivity: $Pr[W=w|X=x]>0$. Under Assumption (1) and (2), $\mu_{w}$ becomes a potential outcome as $\mathbb{E}(Y^{w}|X)=\mathbb{E}(Y|X=x,W=w)$, aiding in estimating $\tau$ and defining the potential outcome $\mu_{w}$.
@@
In this work, we aim to estimate ITE, defined as the difference between potential outcomes $Y^{0}_{i}$ and $Y^{1}_{i}$. However, as observing both outcomes for the same individual $i$ is impossible, often called the fundamental problem of causal inference, previous works focused on estimating the CATE function $\tau(x)$ [3, 4, 5, 6, 7]. When conditioning on the same covariate set, the optimal estimator for the CATE function also serves as the optimal estimator for the ITE function in terms of mean squared error [4]. However, as already stated, CATE approaches predominantly yield point estimates and often overlook quantifying uncertainty in predictions. While the intrinsic value of CATE estimates is undeniable, quantifying uncertainty surrounding these estimates is essential for robust decision-making in high-risk environments [8, 9]. Therefore, it is imperative to address both dimensions of predictive uncertainty, $\widehat{ITE}=\tau(X)+\varepsilon_{a,ITE}(X)+\varepsilon_{e,ITE}(X)$ with the epistemic uncertainty ($\varepsilon_{e,ITE}$) originating from model misspecification and finite samples, and the aleatoric uncertainty ($\varepsilon_{a,ITE}$) arising from the inherent variability of ITE among individuals with the same covariates. Incorporating a comprehensive understanding of these uncertainties is pivotal for advancing the reliability and applicability of heterogeneous treatment effect models.
CATE meta-learners
CATE estimation is frequently performed using CATE meta-learners. CATE meta-learners are frameworks where base learners are utilized to learn and perform CATE estimation [4]. Some examples are the T-learner, S-learner, X-learner [4], DR-learner [6], and R-learner [5]. We can categorize these learners into two groups: one-step plug-in (indirect) learners and two-step (direct) learners. One-step plug-in learners (S- and T-learners) are approaches that use two regression functions where the difference represents the CATE function and thus is estimated indirectly. In contrast, two-step learners (X-, DR-, and R-learners) output the CATE function directly. Three learners are used in this work: the T-, S-, and X-learner. In the binary treatment case, the T-learner fits two different base learners $\mu_{0}$ and $\mu_{1}$ on data without treatment and with treatment, respectively. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}_{1}(X)-\hat{\mu}_{0}(X)$. In contrast, the S-learner fits a single base learner $\mu$ on $[X,W]$ with $W$ the treatment variable. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}(X,W=1)-\hat{\mu}(X,W=0)$. The X-learner starts as a T-learner by first fitting $\mu_{0}$ and $\mu_{1}$, however, afterwards a direct estimate of $\tau$ is fitted with $X$ as input and $W(Y-\hat{\mu}_{0}(X))+(1-W)(\hat{\mu}_{1}(X)-Y)$ as target.
Conformal prediction
Conformal prediction (CP) is a model-agnostic framework that gives an implicit confidence estimate in a prediction by generating prediction sets at a specified significance level $\alpha$ [14]. Notably, this framework allows for (conservatively) valid non-asymptotic confidence predictors under the exchangeability assumption, assuming training/calibration data exchangeability with test data, a slightly less stringent assumption than data independence. The prediction sets in CP are formed by comparing nonconformity scores of examples that quantify how unusual a predicted label is, i.e., these scores measure the disagreement between the prediction and the actual target.
More specifically, we define a prediction interval $\hat{C}(x_{n+1})$, for test object $x_{n+1}\in\mathbf{X}$, by computing following conformity scores $R_{i}^{y}$, based on conformity measure $A$, for each $y\in\mathbb{R}$, using training data $z_{i}=(x_{i},y_{i}),i=1,...,n$:
$R_{i}^{y}=A(z_{i},z_{1:n\setminus i}\cup(x_{n+1},y)),\quad i=1,...,n$ (1)
and
$R_{n+1}^{y}=A((x_{n+1},y),z_{1:n}).$ (2)
The label $y$ is included in prediction interval $\hat{C}(x_{n+1})$ if,
$\frac{|i=1,...,n+1:R_{i}^{y}\geq R_{n+1}^{y}|}{n+1}>\alpha.$ (3)
This procedure, termed full or transductive conformal prediction, is computationally intensive. Therefore, a more applicable variant of full CP, called inductive or split CP (ICP) [20], offers computational efficiency and enables the integration of CP with ML algorithms, such as neural networks and tree-based models. In ICP, the training sequence $(x,y)=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ is portioned into a proper training sequence $\{(x_{1},y_{1}),...,(x_{m},y_{m})\}$ and a calibration sequence $\{(x_{m+1},y_{m+1}),...,(x_{n},y_{n})\}$. We first train a regression model $\hat{y}(x)$ using the proper training sequence. Then, we generate nonconformity scores $R_{i}$ for data points $(x_{i},y_{i})$ where $m<i\leq n$. These scores, which can be calculated using metrics like absolute error $R_{i}=|y_{i}-\hat{y}_{i}|$, are subsequently sorted in descending order, resulting in a ranking: $R^{*}_{1},...,R^{*}_{n-m}$. For a new test object $x_{n+1}$ and point prediction $\hat{y}_{n+1}$, ICP outputs then a prediction interval for a desired target coverage of $1-\alpha$:
$[\hat{y}_{n+1}-R^{*}_{s},\hat{y}_{n+1}+R^{*}_{s}]$ (4)
where $s=\lfloor\alpha(n-m+1)\rfloor$.
Although the conventional ICP yields calibrated prediction intervals, it only offers marginal coverage and constant interval sizes for different covariates. As we want smaller intervals for confident predictions and larger intervals for more "difficult" examples, Mondrian CP [21] and conformalized quantile regression (CQR) [22] extensions of the ICP procedure are proposed. Another adaptation of ICP for establishing more adaptive prediction intervals involves employing normalized nonconformity measures [23]. These normalized measures typically use the product of the absolute error $|y-\hat{f}(x)|$ and the reciprocal of a 1D uncertainty estimate $\frac{1}{\hat{u}(x)}$, expressed as $R=\frac{|y-\hat{f}(x)|}{\hat{u}(x)}$, where $\hat{u}(x)$ may be generated by a distinct model which estimates $|y-\hat{f}(x)|$. Alternatively, $\hat{u}(x)$ can also be obtained using, e.g., the estimated variance from Monte Carlo dropout or an ensemble model [24]. Following a similar procedure as before, we estimate $\hat{u}(x)$ and adjust the predicted output accordingly, resulting in a more adaptive prediction interval:
$[\hat{y}_{n+1}-\hat{u}(x)R^{*}_{s},\hat{y}_{n+1}+\hat{u}(x)R^{*}_{s}]$ (5)
For a more comprehensive exploration of CP, its extensions and applications, we refer the reader to Vovk et al. [14], Angelopoulos and Bates [24], Fontana et al. [25], Manokhin [26].
Conformal predictive systems

### v5

@@ Abstract
Generating probabilistic forecasts of potential outcomes and individual treatment effects (ITE) is essential for risk-aware decision-making in domains such as healthcare, policy, marketing, and finance. We propose two novel methods: the conformal convolution T-learner (CCT) and the conformal Monte Carlo (CMC) meta-learner, that generate full predictive distributions of both potential outcomes and ITEs. Our approaches combine weighted conformal predictive systems with either analytic convolution of potential outcome distributions or Monte Carlo sampling, addressing covariate shift through propensity score weighting. In contrast to other approaches that allow the generation of potential outcome predictive distributions, our approaches are model agnostic, universal, and come with finite-sample guarantees of probabilistic calibration under knowledge of the propensity score. Regarding estimating the ITE distribution, we formally characterize how assumptions about potential outcomes’ noise dependency impact distribution validity and establish universal consistency under independence noise assumptions. Experiments on synthetic and semi-synthetic datasets demonstrate that the proposed methods achieve probabilistically calibrated predictive distributions while maintaining narrow prediction intervals and having performant continuous ranked probability scores. Besides probabilistic forecasting performance, we observe significant efficiency gains for the CCT- and CMC meta-learners compared to other conformal approaches that produce prediction intervals for ITE with coverage guarantees.
@@ 1 Introduction
Modern decision-making requires understanding not just average treatment effects but also complete distributions of potential outcomes. Consider a clinician choosing whether to prescribe a hypertension medication: knowing there is a 60% chance that the drug reduces systolic blood pressure by 15+ mmHg for an individual patient provides much more actionable information than simply knowing it lowers pressure by 12 mmHg on average for a patient with similar characteristics. Similarly, prediction intervals (PI) showing potential effects ranging from 8-16 mmHg offer more insight than point estimates alone, though they still mask important distributional characteristics. Current machine learning approaches often estimate only conditional average treatment effects (CATE)  (Wager and Athey, 2018; Künzel et al., 2019; Nie and Wager, 2021; Kennedy, 2022; Curth and Schaar, 2021), limiting their applicability in uncertain environments, especially in higher-risk settings such as healthcare (Banerji et al., 2023; Feuerriegel et al., 2024). Distribution estimation approaches for individual treatment effect (ITE) address this limitation by characterizing the full distribution of potential outcomes, enabling comparisons between potential outcome distributions. This allows clinicians to assess the expected average benefit of a treatment as well as the complete profile of possible responses for each individual patient. Table 1 compares point, interval, and distributional predictions for a hypertension medication example.
Table 1: Comparison of ITE Estimation Approaches (Hypertension Medication Example). $ITE=Y(1)-Y(0)$ represents individual treatment effect, $\tau$ is a clinical threshold (e.g., 15 mmHg reduction).
Aspect Point Prediction Interval Prediction Distributional Prediction
Output Single estimate: 12 mmHg 90% PI: 8–16 mmHg Full CDF $\hat{F}_{\Psi}(y)$
Threshold Query Cannot quantify $P(ITE\geq 15)$ Bounds contain $15$ but no probability Direct: $P(ITE\geq 15)=0.60$
Risk Assessment Hides all variance Masks bimodality/outliers Reveals probability mass at thresholds
[…]
Possible Theoretical Guarantee Consistent for CATE Marginal coverage Probabilistic and quantile calibration
Reliable distributional estimation for ITEs must address three key challenges: (1e) Confounding robustness to covariate shift via propensity score adjustment; (2) Universal asymptotic convergence to the data-generating ITE distribution; and (3) Finite sample calibration even for small $n$. As can be seen in Table 2, existing distributional approaches struggle mostly with the third challenge to provide calibrated predictive distributions for counterfactuals and ITEs. Additionally, none of the existing solutions is model agnostic. Bayesian methods (BART (Chipman et al., 2010; Hill, 2011), CMGP (Alaa and van der Schaar, 2017), DKLITE (Zhang et al., 2020)) and some deep learning approaches (CEVAE (Louizos et al., 2017)) impose restrictive parametric assumptions limiting universality. While other neural network approaches (GANITE (Yoon et al., 2018), NOFLITE (Vanderschueren et al., 2023), DiffPO (Ma et al., 2024), FCCN (Zhou et al., 2022)) allow for the modeling of more flexible predictive distributions, they often produce poorly calibrated distributions and lack finite-sample guarantees of probabilistic calibration. Conformal prediction (CP) approaches (WCP (Lei and Candès, 2021), CM-learners (Alaa et al., 2023)) advance interval estimation but cannot quantify threshold probabilities or integrate utility functions. These limitations persist because counterfactual inference fundamentally involves unobserved outcomes, making it a challenge to estimate predictive distributions while balancing robustness, calibration, and computational tractability.
Table 2: Comparison of different distributional approaches in the literature.
Approach Methods S-, T-, or Direct- learner Confounding Universal Finite sample Model agnostic
BART (Chipman et al., 2010; Hill, 2011) BART S-learner × × × ×
CEVAE (Louizos et al., 2017) VAE S-learner × × × ×
CMGP (Alaa and van der Schaar, 2017) GP S-learner × × ×
GANITE (Yoon et al., 2018) GAN Direct-learner × × ×
DKLITE (Zhang et al., 2020) Deep kernels S-learner ? × × ×
NOFLITE (Vanderschueren et al., 2023) NFs T or S × × ×
DiffPO (Ma et al., 2024) DM S-learner × × ×
FCCN (Zhou et al., 2022) CN T-learner × ×
CCT-learner (ours) WCPS T-learner
CMC Meta-learner (ours) WCPS T or S
We introduce 3 advances to make distributional regression for causal machine learning more reliable:
1.  Conformalized ITE distribution estimation
We propose a novel model-agnostic framework for deriving nonparametric predictive distributions for counterfactual outcomes and ITEs, based on (weighted) conformal predictive systems (WCPS)(Vovk et al., 2019; Jonkers et al., 2024), a framework that applies (weighted) CP (WCP) (Vovk et al., 2022; Tibshirani et al., 2019) to derive probabilistically calibrated predictive distributions under the (weighted) exchangeability assumption, resulting in two approaches:
  • CCT-learner constructs predictive distributions of ITEs via analytical convolution of propensity-weighted conformal predictive distributions from the potential outcome models, which under certain assumptions guarantee probabilistic calibration of the predictive distribution (Theorem 2).
  • CMC-learner trades finite-sample ITE calibration and some predictive performance for computational efficiency, using Monte Carlo sampling to approximate distributions while still allowing for universal consistency (Theorem 3).
2.  Theoretical guarantees and results
We establish rigorous theoretical foundations for WCPS and our introduced framework:
  • Finite-sample calibration for potential outcome distributions under covariate shift, by proving that WCPS (Jonkers et al., 2024) are guaranteed to be probabilistically calibrated (Theorem 1). This ensures the probabilistic calibration of the potential outcome predictive distributions and allows us to prove that CCT-learners have a finite-sample calibration guarantee under certain assumptions.
  • Existence of universal consistent solution, proving that the CCT- and the CMC meta-learners converge to the data-generating ITE distribution as $n\rightarrow\infty$, assuming the underlying potential outcome models converge (Theorem 3).
  • Noise dependence characterization, formalizing how correlations between noise on potential outcomes affect validity, a critical and often neglected factor in ITE and CATE estimation, especially for uncertainty quantification.
3.  Empirical robustness
Our methods outperform or are on par with existing approaches for prediction interval and predictive distribution estimation across six synthetic and three semi-synthetic datasets, with a total of 88 different settings. The extensive evaluation demonstrates that our methods:
  • Maintain probabilistic calibration even under model misspecification and under propensity score estimation, ensuring reliability in uncertainty quantification across diverse data scenarios.
  • Show superior performance related to continuous ranked probability score (CRPS) across multiple datasets and settings, indicating more accurate predictive distributions than SOTA.
  • Deliver consistent coverage properties in prediction intervals, with empirical coverage rates closely matching nominal coverage levels even in challenging scenarios.
  • Exhibit exceptional flexibility and model-agnosticity, seamlessly integrating with various causal inference models and distribution estimation approaches, highlighting the versatility of our method across different modeling paradigms.
Code and datasets are available at https://github.com/predict-idlab/cct-cmc.
2 Problem definition
@@
In this work, we use the Neyman-Rubin (Splawa-Neyman et al., 1990; Rubin, 2005) potential outcome framework to formulate our problem. We assume access to a sample of observations $\mathcal{D}=\{Z_{i}=(Y_{i},X_{i},W_{i})\}^{n}_{i=1}$, with $Z_{i}\overset{\mathrm{iid}}{\sim}\mathcal{P}$ with $\mathcal{P}$ the resulting joint distribution of the data generation process. Here, $Y_{i}\in\mathcal{Y}$ represents the continuous observed outcome of interest, $X_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ denotes the per-object feature values, and $W_{i}\in\{0,1\}$ is the binary treatment assignment. The joint distribution $\mathcal{P}$ is defined by the following covariate distribution and nuisance functions:
$\pi(x)=\mathbb{P}(W=1|X=x),\quad W\sim\text{Bernoulli}(\pi(x)),\quad X\sim\Lambda$
$Y(w)=\mu_{w}(X)+\varepsilon(w),\quad\mu_{w}(x)=\mathbb{E}(Y|X=x,W=w),\quad Y=Y(0)(1-W)+Y(1)W$
$\tau(x)=\mathbb{E}(Y(1)-Y(0)|X=x),\quad ITE=Y(1)-Y(0)=\tau(X)+\varepsilon_{a,ITE}$
@@
where $\pi(x)$ denotes the propensity score that defines the probability of treatment assignment for an individual, $\mu_{w}(x)$ represents the expected (potential) outcome function of an individual for treatment assignment $w$, $\Lambda$ represents the marginal distribution of $X$, $Y(w)$ represent the potential outcome when an individual receives treatment $w$, $\varepsilon(w)$ are zero-mean random variables independent of $X$ and $W$, $Y$ is the observed outcome, $\tau(x)$ denotes the CATE function, $ITE$ the individual treatment effect (ITE), and $\varepsilon_{a,ITE}$ represent a zero-mean random variable.
Next, we assume that the described data-generating process follows the following assumptions: (1) Consistency: $Y(w_{i})=Y_{i}$; (2) Unconfoundedness: $Y(0),Y(1)\perp W|X$ and (3) Positivity: $Pr[W=w|X=x]>0$. Under Assumption (1) and (2), $\mu_{w}$ becomes a potential outcome as $\mathbb{E}(Y(w)|X)=\mathbb{E}(Y|X=x,W=w)$, aiding in estimating $\tau$ and defining the potential outcome $\mu_{w}$.
@@
In this work, we aim to estimate $ITE_{i}$, defined as the difference between potential outcomes $Y_{i}(1)$ and $Y_{i}(0)$. However, as observing both outcomes for the same individual $i$ is impossible, known as the fundamental problem of causal inference, previous works focused on estimating the CATE function $\tau(x)$ (Wager and Athey, 2018; Künzel et al., 2019; Nie and Wager, 2021; Kennedy, 2022; Curth and Schaar, 2021). When conditioning on the same covariate set, the optimal estimator for the CATE function also serves as the optimal estimator for the ITE in terms of mean squared error (Künzel et al., 2019). However, as already stated, CATE approaches predominantly yield point estimates and often overlook quantifying uncertainty in predictions. While the intrinsic value of CATE estimates is undeniable, quantifying uncertainty surrounding these estimates is essential for robust decision-making in high-risk environments (Banerji et al., 2023; Feuerriegel et al., 2024). Therefore, it is crucial to address both dimensions of predictive uncertainty, $\widehat{ITE}=\hat{\tau}(X)+\varepsilon_{a,ITE}(X)=\tau(X)+\varepsilon_{a,ITE}(X)+\varepsilon_{e,ITE}(X)$ with the epistemic uncertainty ($\varepsilon_{e,ITE}$) originating from model misspecification and finite samples, and the aleatoric uncertainty ($\varepsilon_{a,ITE}=\varepsilon(1)-\varepsilon(0)$) arising from the inherent variability of ITE among individuals with the same covariates. Incorporating a comprehensive understanding of these uncertainties is pivotal for advancing the reliability and applicability of heterogeneous treatment effect models.
3 Conformal predictive systems